OpenStudy (anonymous):

i need to solve for x: x/(a-1)= ax+3

OpenStudy (anonymous):

x/(a-1) - ax=3 x(1/(a-1) -a)=3 x(1-a^2+a)=3 x=3/(a-a^2+1)

OpenStudy (anonymous):